OpenStudy (anonymous):

To determine the calories from physical activity for a moderately active person, what factor should you multiply their BMR by?

OpenStudy (anonymous):

Here are the full equations for finding your BMR. English BMR Formula Women: BMR = 655 + ( 4.35 x weight in pounds ) + ( 4.7 x height in inches ) - ( 4.7 x age in years ) Men: BMR = 66 + ( 6.23 x weight in pounds ) + ( 12.7 x height in inches ) - ( 6.8 x age in year ) Metric BMR Formula Women: BMR = 655 + ( 9.6 x weight in kilos ) + ( 1.8 x height in cm ) - ( 4.7 x age in years ) Men: BMR = 66 + ( 13.7 x weight in kilos ) + ( 5 x height in cm ) - ( 6.8 x age in years )